sqrt(5) * sqrt(10) * sqrt(x^2) =
Found 2 solutions by jim_thompson5910, mananth:
Answer by jim_thompson5910(35256)
Answer by mananth(16946)

sqrt(5) * sqrt(10) * sqrt(x^2) =
= sqrt (5*10) *x
= sqrt(2*5*5)*x
= 5*sqrt2*x
= 5x*sqrt2